OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1)with a slope of 4. y + 3 = 4(x - 1) y - 1 = 4(x - 3) y - 3 = 4(x + 1) y + 1 = 4(x + 3)

OpenStudy (anonymous):

(y-(-1))/(x-(-3))= 4 (y+1)/(x+3) = 4 y+1 = 4(x+3) the last option

OpenStudy (anonymous):

Thank you very much ^_^